OpenStudy (anonymous):

There are two inlet pipes A and B.A fills the tank in 8 hours and B in 12 hours.They are opened at alternate hrs starting with A at 10 am.When will the tank be filled?

OpenStudy (anonymous):

\[\frac{1}{8}*4+\frac{1}{10}*4 = .9\] so its more than 4+4 = 8 hours, but: \[\frac{1}{8}*5+\frac{1}{10}*4 = 1.025\] so its less than 9 hours.

OpenStudy (anonymous):

wait....im retarded

OpenStudy (anonymous):

gotta flip those numbers

OpenStudy (anonymous):

The options are :- 1) 21/2 2) 10 3) 19/2 4) 28/3

OpenStudy (anonymous):

no wait...ive confused myself <.<

OpenStudy (anonymous):

nvm, i stand by what i said in my first post lol

OpenStudy (anonymous):

Is it c) 19/2 hrs

OpenStudy (anonymous):

i think its 8 hrs and 48 mins.

OpenStudy (anonymous):

it is 19/2

OpenStudy (anonymous):

opened at alternate hours got to add it up

OpenStudy (anonymous):

\[\frac{1}{8}+\frac{1}{12}+...=1\]

OpenStudy (anonymous):

i guess im not seeing it correctly then? i thought that what i was doing.

OpenStudy (anonymous):

OH, i saw a 10, but its 12 in the question.

OpenStudy (anonymous):

i must have looked at the 10 AM part >.< bah.

OpenStudy (anonymous):

takes 9 hours and and a little

OpenStudy (anonymous):

oh right. nine and a half

OpenStudy (anonymous):

how did you get that so fast? or did you do it slow?

OpenStudy (anonymous):

tank filled at 7:30

OpenStudy (anonymous):

it's obvious that it takes more than (5 hrs from A + 4 hrs from B) and less than (5 hrs from A + 5 hrs from B) since 5/8 + 4/12 = 0.9583 5/8 + 5/12 = 25/24 > 1 since A starts first 5 hrs of A have to be completed and (4+ ?) hrs [ say x ]of B therefore 5/8 + x/12 = 1 => x = 9/2 so total time = 5 hrs of A + 9/2 hrs of B = 19/2 hrs

OpenStudy (anonymous):

ok you actually computed but with thought. nice work